# booktabs, siunitx

### v1

@@ Abstract
Over the last few years, primordial black holes (PBHs) have emerged as a strong candidate for cold dark matter. A significant number of PBHs are produced when the strength of the primordial scalar power spectrum is enhanced on small scales (compared to the COBE normalized values on large scales). Such primordial spectra also inevitably lead to strong amplification of the scalar-induced, secondary gravitational waves (GWs) at higher frequencies. The recent detection of the stochastic gravitational wave background (SGWB) by the pulsar timing arrays (PTAs) has opened up the possibility of directly probing the very early universe. Different studies have shown that, when PBHs are assumed to have been formed during the epoch of radiation domination, the mechanism for the amplification of the scalar-induced GWs that is required to explain the PTA data can overproduce the PBHs over some ranges of masses. In this work, we assume a specific functional form for the primordial scalar power spectrum and examine the production of PBHs and the scalar-induced secondary GWs during the phase of reheating, which precedes the standard epoch of radiation domination. Specifically, we account for the uncertainties in the conditions for the formation of PBHs and ensure that the extent of PBHs produced remains within the observational bounds. We find that the scalar-induced SGWB generated during a phase of reheating with a steeper equation of state (than that of radiation) fit the NANOGrav 15-year data with a stronger Bayesian evidence than the astrophysical scenario involving GWs produced by merging supermassive binary black holes.
@@ 1 Introduction
During the last few decades, the observations of the cosmic microwave background (CMB) has provided the most effective tool to probe the early universe prior to the surface of the last scattering. The increasingly accurate measurements of the anisotropies in the CMB by COBE [1, 2], WMAP [3, 4, 5, 6] and Planck [7, 8, 9] have defined the era of precision cosmology. These observations have led to strong constraints on the parameters that characterize the primordial perturbations such as the scalar amplitude $A_{{}_{\mathrm{S}}}$, the scalar spectral index $n_{{}_{\mathrm{S}}}$, the tensor-to-scalar ratio $r$, and the scalar non-Gaussianity parameter $f_{{}_{\mathrm{NL}}}$ [10, 11]. However, the CMB constrains the primordial perturbations only on large scales (i.e. wave numbers corresponding to $10^{-4}<k<1\,\mathrm{Mpc}^{-1}$), and the corresponding bounds on smaller scales are considerably weaker.
@@
Over the past decade, the detection of gravitational waves (GWs) by the Ligo-Virgo-Kagra (LVK) collaboration (in this context, see, for instance, Refs. [12, 13, 14, 15, 16, 17]) has led to a paradigm shift in the approach to primordial cosmology. It is now widely recognized that the GWs provide a unique window to probe the primordial physics (see, for example, the recent reviews [18, 19, 20, 21]). The GWs observed by the LVK detectors were of astrophysical origin, and they were produced by the mergers of binary black holes or neutron stars. As we shall discuss below, there exist a variety of phenomena that can generate GWs in the primordial universe and, importantly, the GWs produced in the early universe will be stochastic in nature. Moreover, because they interact weakly, GWs carry information about the earliest stages of our universe and, hence, they probe primordial physics more effectively than the CMB, in particular, on small scales. But, despite a dedicated effort, the LVK collaboration did not detect a stochastic GW background (SWGB) over the frequency range of, say, $1<f<10^{2}\,\mathrm{Hz}$, that its detectors are sensitive to [22]. However, a few months ago, this barrier was overcome when the Pulsar Timing Arrays (PTAs)—such as NANOGrav [23, 24], EPTA (including the data from InPTA) [25, 26], PPTA [27, 28], and CPTA [29]—jointly reported the first detection of a SGWB over the frequency range of $10^{-9}<f<10^{-6}\,\mathrm{Hz}$. Specifically, it was found that the SGWB generated by the merging of supermassive black hole binaries (SMBHBs) do not statistically explain the observed data in an efficient manner [30]. In fact, the NANOGrav 15-year data favored (with a particularly strong Bayesian evidence) primordial origins for the detected SGWB signal [30, 24, 23]. This has immediately fuelled the search for several different channels for the cosmological generation of the SGWB. These include the SGWB from: (i) models that lead to enhanced formation of PBHs [31, 32, 33, 34, 35, 36, 37, 38, 39, 40], (ii) domination of ultra-low mass PBHs [41], (iii) mergers due to PBHs [42, 43], (iv) phase transitions in the early universe [44, 45, 46, 47, 48, 49, 50, 51], (v) collapse of cosmic strings and domain walls [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62], (vi) models of axion inflation [36, 63, 64], (vii) blue tilted primary tensor power spectra [65, 66, 67, 68], and (viii) primordial magnetic fields [69].
@@
In this work, we focus on SGWB generated during the formation of PBHs. This is one of the most studied explanations for the PTA data. Nevertheless, we believe that, in the scenario, there are some caveats, such as the overproduction of PBHs, which require further investigation. PBHs are an attractive candidate for dark matter because of their cold, non-interacting nature [70, 71, 72, 73]. Contrary to astrophysical BHs, which are typically of the order of solar mass, PBHs can have masses ranging from a few grams to billions of solar masses. PBHs can also play a pivotal role in shaping our understanding of the early universe. While they are constrained in a variety of manners over different mass ranges, there still remains a window over the asteroid mass ranges where the PBHs can account for the dark matter completely [74]. There are a number of potential mechanisms through which PBHs can form, including bubble collisions [75, 76, 77], collapse of domain walls [78, 79, 80, 81, 82, 83], collapse of cosmic strings [84, 85, 86, 87, 88, 89], electroweak phase transition [90], first-order phase transitions [91, 92], and formation of other topological defects [93, 94]. However, in this work, we shall be interested in the scenario wherein PBHs are formed due to the collapse of the enhanced density perturbations that originated during inflation [95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105].
@@
Inflation corresponds to the earliest phase of accelerated expansion of our universe. Though it was initially proposed to resolve the horizon problem, the success of inflation lies in the fact that it also provides a mechanism for the creation of the primordial scalar and tensor perturbations (in this regard, see the reviews [106, 107, 108, 109, 110, 111, 112]). While the inflationary scalar perturbations contribute to the density fluctuations and result in the emergence of structure in the universe, the tensor perturbations correspond to GWs. These perturbations are eventually responsible for the anisotropies observed in the CMB and the distribution of the large-scale structure that we see around us today. The observations of the anisotropies in the CMB by missions such as Planck [113, 10] and BICEP/Keck [114] allow us to strongly constrain the primordial scalar power spectrum around the pivot scale of $k_{\ast}=0.05\,\mathrm{Mpc}^{-1}$. But, as we mentioned, the constraints on the primordial scalar power spectrum are considerably weaker over smaller scales, say, corresponding to wave numbers $k\gg 10^{5}\,\mathrm{Mpc}^{-1}$. Due to this reason, the scalar power spectrum can, in principle, be considerably amplified on small scales (when compared to the COBE normalized values on large scales), leading to the formation of a significant number of PBHs when the concerned wave numbers re-enter the Hubble radius after inflation. Such power spectra with enhanced power on small scales can be generated during inflation with the aid of potentials that contain either a point of inflection [115, 97, 100, 102, 116, 117, 118, 119, 120] or a small bump/dip [121]. They can also be generated in inflationary models involving more than one field [122, 123, 124], due to resonant amplification during inflation [125, 126] or from squeezed initial states [127]. At the first order in perturbation theory, the scalar and tensor perturbations are decoupled and hence evolve independently. However, at the higher order in the perturbations, one form of perturbation can act as a source for the other. As a result, when the amplitude of the scalar perturbations at the first order is enhanced over small scales to produce a significant number of PBHs, they inevitably induce tensor perturbations of considerable strengths at the second order. In such scenarios, the PBHs form almost instantaneously when the wave numbers with significant scalar power re-enter the Hubble radius and, in the process, also generate a SGWB of amplitudes that are comparable to the sensitivities of the ongoing and forthcoming GW observatories.
@@
The formation of PBHs and the generation of scalar-induced, secondary GWs during the epoch of radiation domination have been discussed in the literature in good detail, particularly in the context of the PTA data. The main caveat in this scenario remains the fine-tuning and the possible overproduction of PBHs in a certain range of masses [31]. This issue has motivated further studies wherein the formation of PBHs has been considered during non-standard histories post inflation [128, 129, 40, 130, 131, 132]. Reheating is the phase that immediately follows inflation. During this period, the inflaton oscillates at the bottom of the inflationary potential, and, with the aid of suitable couplings, the energy from the inflaton is transferred to the fields that constitute the standard model of particle physics. In principle, the PBHs can form during the phase of reheating and, in such a scenario, the parameters such as the effective equation-of-state (EoS) during reheating and the reheating temperature will affect the formation mechanism as well as the number of PBHs produced [133]. An epoch of reheating with a general EoS can also significantly alter the strength and the spectral shape of the secondary GWs [134, 135, 20]. In this study, we focus on the formation of PBHs during the phase of reheating and explore the manner in which such a scenario can explain the observed PTA data. We take into account the uncertainties in the conditions for the collapse of the PBHs as well as the existing bounds on the abundance of PBHs in the relevant range of masses. We assume an analytical expression for the inflationary scalar power spectrum, which is in the form of a broken power law. Such a spectral shape is motivated by the forms of the scalar power spectra that typically arise when one considers models of inflation that permit a brief epoch of ultra slow roll [136, 120]. We scan our parameter space for an optimal fit to the NANOGrav 15-year data, while staying strictly within the observational bounds on the abundance of the PBHs. We perform a Bayesian analysis and compare our scenario with the possible astrophysical explanation for the PTA results through the mergers of SMBHBs. We find that when we allow room for uncertainties in the conditions for the collapse of PBHs, our model is strongly preferred over the scenario involving SMBHBs.
@@ 2.1 A brief overview of the dynamics during reheating
Due to the transfer of energy from the inflaton, the dimensionless energy density of radiation, viz. $\rho_{{}_{\mathrm{R}}}/(\rho_{{}_{\mathrm{R}}}+\rho_{\phi})$, keeps growing until it approaches unity, marking the end of reheating at a given conformal time, say, $\eta_{\mathrm{re}}$. It is often found that, throughout the entire period, except towards the end, the reheating phase is governed by the inflaton EoS $w_{\phi}$. Hence, we shall assume that the EoS is constant throughout the reheating phase at a given value $w_{\mathrm{re}}=w_{\phi}$. This allows us to describe the reheating phase in terms of only two parameters: the EoS during reheating $w_{\mathrm{re}}$ and the reheating temperature $T_{\mathrm{re}}$, which is the temperature of the radiation when reheating is complete.
@@ 2.2 Effects of reheating on the formation of PBHs
In this section, we shall discuss the formation of PBHs due to the enhancement in power on small scales in the spectrum of curvature perturbations generated during inflation. In the simple picture wherein reheating occurs instantaneously after inflation, all the small-scale modes re-enter the Hubble radius during the epoch of radiation domination. When scalar perturbations with significant amplitudes re-enter the Hubble radius, the overdense region collapses almost instantaneously to form the PBHs. The mass of the PBHs thus produced turns out to be nearly equivalent to the mass within the Hubble radius at the time of re-entry. When there occurs a phase of reheating, some of the small-scale modes re-enter the Hubble radius during reheating. Even in such a situation, as the reheating dynamics can be described by an effective EoS, the mechanism for the formation of the PBHs proves to be the same. However, apart from the amplitude of the scalar power spectrum, the efficiency of the formation of PBHs during reheating depends on the EoS. We shall focus on the collapse of the overdense regions due to the large scalar perturbations generated during inflation, which ensures instantaneous collapse. We should mention that the location of the peak in the inflationary scalar power spectrum determines the time of re-entry of that scale, and thus the time of PBH formation. Late-forming PBHs correspond to peaks at smaller wave numbers and thus lead to PBHs of higher masses. PBHs with mass $M<10^{15}\,\mathrm{g}$ would have evaporated by today due to Hawking radiation. In this work, we shall be interested in PBHs with mass $M>10^{15}\,\mathrm{g}$, which can contribute to the cold dark matter.
In our analysis, we shall not consider any specific inflationary potential for generating the scalar power spectrum. Rather, we shall consider an analytical, scalar power spectrum with a broken power law form. On large scales, we shall assume that the power spectrum is nearly scale invariant, as is required to fit the CMB data. On small scales, we shall assume that the power spectrum rises as $k^{4}$ as it approaches the peak and, beyond the peak, it falls as $k^{n_{0}}$ with $n_{0}<0$. We should point out that such power spectra arise in single-field models of inflation that permit a brief period of ultra slow roll [136]. The complete form of the scalar power spectrum that we shall work with is given by
$\mathcal{P}_{\mathcal{R}}(k)=A_{{}_{\mathrm{S}}}\left(\frac{k}{k_{\ast}}\right)^{n_{{}_{\mathrm{S}}}-1}+A_{0}\left\{\begin{aligned} &\left(\frac{k}{k_{\mathrm{peak}}}\right)^{4}\hskip 1.00006pt&k\leq k_{\mathrm{peak}}\\ &\left(\frac{k}{k_{\mathrm{peak}}}\right)^{n_{0}}\hskip 1.00006pt&k\geq k_{\mathrm{peak}}\end{aligned}\right.,$ (2.8)
@@
If $\sigma_{\delta}(R)$ represents the variance in the density fluctuations in space, that is arrived at by smoothing over a radius $R$ with the help of the window function $W(k\,R)$, then the quantity is given in terms of the power spectrum $\mathcal{P}_{\delta}(k)$ through the relation
@@
Let $M_{{}_{\mathrm{H}}}$ be the total mass within the Hubble radius $H^{-1}$ at a given time, and let $M_{\mathrm{eq}}$ be the mass within the Hubble radius at the time of radiation-matter equality. The mass $M$ of the PBHs formed due to collapse is related to the quantity $M_{{}_{\mathrm{H}}}$ at the time of formation as follows:
@@
where $\rho_{\rm total}$ denotes the total energy density of the background and $\gamma$ is the efficiency of collapse given by $\gamma=w_{\mathrm{re}}^{3/2}$ [152]. For a given reheating temperature $T_{\mathrm{re}}$, we can relate the mass $M$ of the PBHs at the time of their formation to the quantity $M_{\mathrm{eq}}$ as
@@
It is evident from our analysis that the mass $M$ of the PBHs and the quantity $f_{{}_{\mathrm{PBH}}}(M)$ not only depend on the inflationary scalar power spectrum but also on the parameters describing the epoch of reheating. However, for any fixed shape of the power spectra, certain behavior, such as $f_{{}_{\mathrm{PBH}}}(M)\propto{T_{\mathrm{re}}}^{\frac{1-3\,w_{\mathrm{re}}}{1+w_{\mathrm{re}}}}\,{M}^{-\frac{2\,w_{\mathrm{re}}}{1+w_{\mathrm{re}}}}$ remains valid irrespective of the shape, while the individual values of PBH mass fraction will naturally depend on the shape of the inflationary spectrum. We should also point out that, for the power spectrum (2.8) we have considered, the $f_{{}_{\mathrm{PBH}}}(M)$ proves to be nearly monochromatic in nature.

### v2

# Constraining the history of reheating with the NANOGrav 15-year data

@@ Abstract
Over the last few years, primordial black holes (PBHs) have emerged as a strong candidate for cold dark matter. A significant number of PBHs are produced when the strength of the primordial scalar power spectrum is enhanced on small scales (compared to the COBE normalized values on large scales). Such primordial spectra also inevitably lead to strong amplification of the scalar-induced, secondary gravitational waves (GWs) at higher frequencies. The recent detection of the stochastic GW background (SGWB) by the pulsar timing arrays (PTAs) has opened up the possibility of directly probing the very early universe. Different studies have shown that, when PBHs are assumed to have been formed during the epoch of radiation domination, the mechanism for the amplification of the scalar-induced GWs that is required to explain the PTA data can overproduce the PBHs over some ranges of masses. In this work, we assume a specific functional form for the primordial scalar power spectrum and examine the production of PBHs and the scalar-induced secondary GWs during the phase of reheating, which precedes the standard epoch of radiation domination. Specifically, we account for the uncertainties in the conditions for the formation of PBHs and ensure that the extent of PBHs produced remains within the observational bounds. We find that the scalar-induced SGWB generated during a phase of reheating with a steeper equation of state (than that of radiation) fit the NANOGrav 15-year data with a stronger Bayesian evidence than the astrophysical scenario involving GWs produced by merging supermassive binary black holes.
@@ 1 Introduction
During the last few decades, the observations of the cosmic microwave background (CMB) have provided the most effective tool to probe the early universe prior to the surface of the last scattering. The increasingly accurate measurements of the anisotropies in the CMB by COBE [1, 2], WMAP [3, 4, 5, 6] and Planck [7, 8, 9] have defined the era of precision cosmology. These observations have led to strong constraints on the parameters that characterize the primordial perturbations such as the scalar amplitude $A_{{}_{\mathrm{S}}}$, the scalar spectral index $n_{{}_{\mathrm{S}}}$, the tensor-to-scalar ratio $r$, and the scalar non-Gaussianity parameter $f_{{}_{\mathrm{NL}}}$ [10, 11]. However, the CMB constrains the primordial perturbations only on large scales (i.e. wave numbers corresponding to $10^{-4}<k<1\,\mathrm{Mpc}^{-1}$), and the corresponding bounds on smaller scales are considerably weaker.
@@
Over the past decade, the detection of gravitational waves (GWs) by the Ligo-Virgo-Kagra (LVK) collaboration (in this context, see, for instance, Refs. [12, 13, 14, 15, 16, 17]) has led to a paradigm shift in the approach to primordial cosmology. It is now widely recognized that GWs provide a unique window to probe the physics operating in the early universe (see, for example, the recent reviews [18, 19, 20, 21]). The GWs observed by the LVK detectors were of astrophysical origin, and they were produced by the mergers of binary black holes (BHs) or neutron stars. As we shall discuss below, there exist a variety of phenomena that can generate GWs in the primordial universe and, importantly, the GWs produced in the early universe will be stochastic in nature. Moreover, because they interact weakly, GWs carry information about the earliest stages of our universe and, hence, they probe primordial physics more effectively than the CMB, in particular, on small scales. But, despite a dedicated effort, the LVK collaboration did not detect a stochastic GW background (SWGB) over the frequency range of, say, $1<f<10^{2}\,\mathrm{Hz}$, that its detectors are sensitive to [22]. However, a few months ago, this barrier was overcome when the Pulsar Timing Arrays (PTAs)—such as NANOGrav [23, 24], EPTA (including the data from InPTA) [25, 26], PPTA [27, 28], and CPTA [29]—jointly reported the first detection of an SGWB over the frequency range of $10^{-9}<f<10^{-6}\,\mathrm{Hz}$. Specifically, it was found that the SGWB generated by the merging of supermassive BH binaries (SMBHBs) do not statistically explain the observed data in an efficient manner [30]. In fact, the NANOGrav 15-year data favored (with a particularly strong Bayesian evidence) primordial origins for the detected SGWB signal [30, 24, 23]. This has immediately fuelled the search for several different channels for the cosmological generation of the SGWB. These include the SGWB from: (i) models that lead to enhanced formation of primordial BHs (PBHs) [31, 32, 33, 34, 35, 36, 37, 38, 39, 40], (ii) domination of ultra-low mass PBHs [41], (iii) mergers due to PBHs [42, 43], (iv) phase transitions in the early universe [44, 45, 46, 47, 48, 49, 50, 51], (v) collapse of cosmic strings and domain walls [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62], (vi) models of axion inflation [36, 63, 64], (vii) blue tilted primary tensor power spectra [65, 66, 67, 68], (viii) primordial magnetic fields [69] and (ix) other possible scenarios [70, 71, 72].
@@
In this work, we focus on SGWB generated during the formation of PBHs. This is one of the most studied explanations for the PTA data. Nevertheless, we believe that, in the scenario, there are some caveats, such as the overproduction of PBHs, which require further investigation. PBHs are an attractive candidate for dark matter because of their cold, non-interacting nature [73, 74, 75, 76]. Contrary to astrophysical BHs, which are typically of the order of solar mass, PBHs can have masses ranging from a few grams to billions of solar masses. PBHs can also play a pivotal role in shaping our understanding of the early universe. While they are constrained in a variety of manner over different mass ranges, there still remains a window over the asteroid mass range where the PBHs can account for the dark matter completely [77]. There are a number of potential mechanisms through which PBHs can form, including bubble collisions [78, 79, 80], collapse of domain walls [81, 82, 83, 84, 85, 86], collapse of cosmic strings [87, 88, 89, 90, 91, 92], electroweak phase transition [93], first-order phase transitions [94, 95], and formation of other topological defects [96, 97]. However, in this work, we shall be interested in the scenario wherein PBHs are formed due to the collapse of the enhanced density perturbations that originated during inflation [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108].
@@
Inflation corresponds to the earliest phase of accelerated expansion of our universe. Though it was initially proposed to resolve the horizon problem, the success of inflation lies in the fact that it also provides a mechanism for the creation of the primordial scalar and tensor perturbations (in this regard, see the reviews [109, 110, 111, 112, 113, 114, 115]). While the inflationary scalar perturbations contribute to the density fluctuations and result in the emergence of structure in the universe, the tensor perturbations correspond to GWs. These perturbations are eventually responsible for the anisotropies observed in the CMB and the distribution of the large-scale structure that we see around us today. The observations of the anisotropies in the CMB by missions such as Planck [116, 10] and BICEP/Keck [117] allow us to strongly constrain the primordial scalar power spectrum around the pivot scale of $k_{\ast}=0.05\,\mathrm{Mpc}^{-1}$. But, as we mentioned, the constraints on the primordial scalar power spectrum are considerably weaker over smaller scales, say, corresponding to wave numbers $k\gg 10^{5}\,\mathrm{Mpc}^{-1}$. Due to this reason, the scalar power spectrum can, in principle, be considerably amplified on small scales (when compared to the COBE normalized values on large scales), leading to the formation of a significant number of PBHs when the concerned wave numbers re-enter the Hubble radius after inflation. Such power spectra with enhanced power on small scales can be generated during inflation with the aid of potentials that contain either a point of inflection [118, 100, 103, 105, 119, 120, 121, 122, 123] or a small bump/dip [124]. They can also be generated in inflationary models involving more than one field [125, 126, 127, 128, 129], due to resonant amplification during inflation [130, 131] or from squeezed initial states [132]. At the first order in perturbation theory, the scalar and tensor perturbations are decoupled and hence evolve independently. However, at the higher order in the perturbations, one form of perturbation can act as a source for the other. As a result, when the amplitude of the scalar perturbations at the first order is enhanced over small scales to produce a significant number of PBHs, they inevitably induce tensor perturbations of considerable strengths at the second order. In such scenarios, the PBHs form almost instantaneously when the wave numbers with significant scalar power re-enter the Hubble radius and, in the process, also generate a SGWB of amplitudes that are comparable to the sensitivities of the ongoing and forthcoming GW observatories.
@@
The formation of PBHs and the generation of scalar-induced, secondary GWs during the epoch of radiation domination have been discussed in the literature in good detail, particularly in the context of the PTA data. The main caveat in this scenario remains the fine-tuning and the possible overproduction of PBHs in a certain range of masses [31]. This issue has motivated further studies wherein the formation of PBHs has been considered during non-standard histories post inflation [133, 134, 40, 135, 136, 137]. Reheating is the phase that immediately follows inflation. During this period, the inflaton oscillates at the bottom of the inflationary potential, and, with the aid of suitable couplings, the energy from the inflaton is transferred to the fields that constitute the standard model of particle physics. In principle, the PBHs can form during the phase of reheating and, in such a scenario, the parameters such as the effective equation-of-state (EoS) during reheating and the reheating temperature will affect the formation mechanism as well as the number of PBHs produced [138]. An epoch of reheating with a general EoS can also significantly alter the strength and the spectral shape of the secondary GWs [139, 140, 20]. In this study, we focus on the formation of PBHs during the phase of reheating and explore the manner in which such a scenario can explain the observed PTA data. We take into account the uncertainties in the conditions for the collapse of the PBHs as well as the existing bounds on the abundance of PBHs in the relevant range of masses. We assume an analytical expression for the inflationary scalar power spectrum, which is in the form of a broken power law. Such a spectral shape is motivated by the forms of the scalar power spectra that typically arise when one considers models of inflation that permit a brief epoch of ultra slow roll [141, 123]. We scan our parameter space for an optimal fit to the NANOGrav 15-year data, while staying strictly within the observational bounds on the abundance of the PBHs. We perform a Bayesian analysis and compare our scenario with the possible astrophysical explanation for the PTA results through the mergers of SMBHBs. We find that, when we allow room for uncertainties in the conditions for the collapse of PBHs, our model is strongly preferred over the scenario involving SMBHBs.
@@ 2.1 A brief overview of the dynamics during reheating
Due to the transfer of energy from the inflaton, the dimensionless energy density of radiation, viz. $\rho_{{}_{\mathrm{R}}}/(\rho_{{}_{\mathrm{R}}}+\rho_{\phi})$, keeps growing until it approaches unity, marking the end of reheating at a given conformal time, say, $\eta_{\mathrm{re}}$. It is often found that, throughout the entire period, except towards the end, the reheating phase is governed by the inflaton EoS $w_{\phi}$. Hence, we shall assume that the EoS is constant throughout the reheating phase at a given value $w_{\mathrm{re}}=w_{\phi}$. This allows us to describe the reheating phase in terms of only two parameters: the EoS parameter during reheating $w_{\mathrm{re}}$ and the reheating temperature $T_{\mathrm{re}}$, which is the temperature of the radiation when reheating is complete.
@@ 2.2 Effects of reheating on the formation of PBHs
In this section, we shall discuss the formation of PBHs due to the enhancement in power on small scales in the spectrum of curvature perturbations generated during inflation. In the simple picture wherein reheating occurs instantaneously after inflation, all the small-scale modes re-enter the Hubble radius during the epoch of radiation domination. When scalar perturbations with significant amplitudes re-enter the Hubble radius, the overdense region collapses almost instantaneously to form PBHs. The mass of the PBHs thus produced turns out to be nearly equivalent to the mass within the Hubble radius at the time of re-entry. When there occurs a phase of reheating, some of the small-scale modes re-enter the Hubble radius during reheating. Even in such a situation, as the reheating dynamics can be described by an effective EoS, the mechanism for the formation of the PBHs proves to be the same. However, apart from the amplitude of the scalar power spectrum, the efficiency of the formation of PBHs during reheating depends on the EoS. We shall focus on the collapse of the overdense regions due to the large scalar perturbations generated during inflation, which ensures instantaneous collapse. We should mention that the location of the peak in the inflationary scalar power spectrum determines the time of re-entry of that scale, and thus the time of PBH formation. Late-forming PBHs correspond to peaks at smaller wave numbers and thus lead to PBHs of higher masses. PBHs with mass $M<10^{15}\,\mathrm{g}$ would have evaporated by today due to Hawking radiation. In this work, we shall be interested in PBHs with mass $M>10^{15}\,\mathrm{g}$, which can contribute to the cold dark matter.
In our analysis, we shall not consider any specific inflationary potential for generating the scalar power spectrum. Rather, we shall consider an analytical scalar power spectrum with a broken power law form. On large scales, we shall assume that the power spectrum is nearly scale invariant, as is required to fit the CMB data. On small scales, we shall assume that the power spectrum rises as $k^{4}$ as it approaches the peak and, beyond the peak, it falls as $k^{n_{0}}$ with $n_{0}<0$. We should point out that such power spectra arise in single-field models of inflation that permit a brief period of ultra slow roll [141]. The complete form of the scalar power spectrum that we shall work with is given by
$\mathcal{P}_{\mathcal{R}}(k)=A_{{}_{\mathrm{S}}}\left(\frac{k}{k_{\ast}}\right)^{n_{{}_{\mathrm{S}}}-1}+A_{0}\left\{\begin{aligned} &\left(\frac{k}{k_{\mathrm{peak}}}\right)^{4}\hskip 1.00006pt&k\leq k_{\mathrm{peak}},\\ &\left(\frac{k}{k_{\mathrm{peak}}}\right)^{n_{0}}\hskip 1.00006pt&k\geq k_{\mathrm{peak}},\end{aligned}\right.$ (2.8)
@@
To calculate the abundance of PBHs, the two most popular methods that are often used in the literature are the Press-Schechter formalism [156] and peak theory [157]. In general, the results from these two established methods for calculating the probability of collapse to form PBHs do not converge, even when all the other conditions and parameters are fixed (in this context, see Refs. [158, 159, 160, 122]). Also, there have been recent efforts from different groups to understand the efficiency of collapse leading to the formation of PBHs with the aid of more involved and accurate numerical frameworks using the so-called compaction function methods (see, for example, Refs. [161, 162, 163, 164, 165, 166, 167]). These efforts indicate that small differences in the details of implementation lead to rather different final results. In this work, we shall follow the simple Press-Schechter formalism which considers a Gaussian random field for the initial density fluctuations and a constant critical density contrast for the formation of PBHs.
@@
If $\sigma_{\delta}(R)$ represents the variance in the density fluctuations in space, which is arrived at by smoothing over a radius $R$ with the help of the window function $W(k\,R)$, then the quantity is given in terms of the power spectrum $\mathcal{P}_{\delta}(k)$ through the relation
@@
Let $M_{{}_{\mathrm{H}}}$ be the total mass within the Hubble radius $H^{-1}$ at a given time, and let $M_{\mathrm{eq}}$ be the mass within the Hubble radius at the time of radiation-matter equality. The mass $M$ of the PBHs formed due to collapse is related to the quantity $M_{{}_{\mathrm{H}}}$ at the time of formation as follows [169]:
@@
where $\rho_{\rm total}$ denotes the total energy density of the background and $\gamma$ is the efficiency of collapse given by $\gamma=w_{\mathrm{re}}^{3/2}$  [^1] At this point, we should mention that there exists a more accurate method to estimate the mass $M$ of the PBHs, with the mass being sensitive to the EoS parameter, say, $w_{\mathrm{re}}$, of the cosmological background fluid as well as the profile of the scalar fluctuations [170, 171, 172, 173, 174, 175, 176, 177]. In the study of critical phenomena in gravitational collapse, numerical simulations have demonstrated that the mass of PBHs obey a scaling relation when the amplitude of the perturbation is close to the critical value (i.e. when $\delta_{\rm m}-\delta_{\rm c}\leq 10^{-2}$, where $\delta_{\rm m}$ is the initial amplitude of the perturbation). According to the scaling relation, the mass of the PBHs is given by $M=\kappa\,M_{{}_{\rm H}}\,\left(\delta_{\rm m}-\delta_{\mathrm{c}}\right)^{\gamma_{\rm f}},$ (2.13) where the exponent $\gamma_{\rm f}$ depends on the EoS parameter $w_{\mathrm{re}}$, while the overall constant $\kappa$ and the critical density $\delta_{\mathrm{c}}$ depend both on $w_{\mathrm{re}}$ and the shape of the scalar power spectrum (in this context, see Refs. [178, 170, 174]; for a discussion on the dependence of $\gamma_{\rm f}$ on $w_{\mathrm{re}}$, see Ref.[179], Tab. 1 and Ref. [180], Fig. 7). However, we find that analytical estimates of $\gamma_{\mathrm{f}}$ and $\kappa$, which would have been helpful, are yet to be arrived at in the literature.. For a given reheating temperature $T_{\mathrm{re}}$, we can relate the mass $M$ of the PBHs at the time of their formation to the quantity $M_{\mathrm{eq}}$ as
@@
It is evident from our analysis that the mass $M$ of the PBHs and the quantity $f_{{}_{\mathrm{PBH}}}(M)$ not only depend on the inflationary scalar power spectrum but also on the parameters describing the epoch of reheating. However, certain behavior, such as $f_{{}_{\mathrm{PBH}}}(M)\propto{T_{\mathrm{re}}}^{\frac{1-3\,w_{\mathrm{re}}}{1+w_{\mathrm{re}}}}\,{M}^{-\frac{2\,w_{\mathrm{re}}}{1+w_{\mathrm{re}}}}$ remains valid irrespective of the shape of the power spectrum, whereas the actual value of the PBH mass fraction naturally depends on all the parameters involved. We should also point out that, for the power spectrum (2.8) we have considered, the $f_{{}_{\mathrm{PBH}}}(M)$ proves to be nearly monochromatic in nature.
There is another point that we need to clarify at this stage of our discussion. Note that the pre-factor $(1+w_{\mathrm{re}})/(5+3\,w_{\mathrm{re}})$ in $\mathcal{P}_{\delta}(k)$ and $\delta_{\mathrm{c}}^{\mathrm{an}}$ [see Eqs. (2.10) and (2.15)], which originates from the Poisson equation, cancel each other on substituting them in the expression (2.9) for $\beta(M)$. We find that the term $\sin^{2}[\pi\,\sqrt{w_{\mathrm{re}}}/(1+3\,w_{\mathrm{re}})]$ that appears in the expression (2.15) for $\delta_{\mathrm{c}}^{\mathrm{an}}$ has a global maximum at $w_{\mathrm{re}}=1/3$. Since the maximum of $\delta_{\mathrm{c}}^{\mathrm{an}}$ occurs at $w_{\mathrm{re}}=1/3$, the abundance of PBHs has a minimum at the value. It is due to this reason that, at the time of radiation-matter equality, the quantity $f_{{}_{\mathrm{PBH}}}(M)$ exhibits a similar dependence on $w_{\mathrm{re}}$, as we have shown in the right panel of Fig. 1. However, we should add that such a dependence of $\beta(M)$ on $w_{\mathrm{re}}$ follows from the formalism we have adopted to estimate the number of PBHs formed [187]. Such a feature may not arise if one follows a different approach to calculate the abundance of PBHs (in this regard, see, for example, Ref. [137]).